OpenStudy (anonymous):

if x=3a+7 and y=6+a, then what expresses y in terms of x?

OpenStudy (anonymous):

x= 3a+7 therefore 3a= x-7 so a= (x-7)/3 substituted 'a' in y equation. y= (x-7)/3 +6